Calculate molecular mass of glucose (C6H12O6) molecule.

Grade:12

3 Answers

Anjali Ahuja
9 years ago

Molar mass of glucose (C6H12O6) = 6*C + 12*H + 6*O
= 6*12 + 12*1 + 6*16
=180 g /mol

!!!!!!!!!!!!!!!!!!!!!.......................................................!!!!!!!!!!!!!!!!!!!!!!!
Anjali Ahuja
9 years ago

Molar mass of glucose (C6H12O6) = 6*C + 12*H + 6*O
= 6*12 + 12*1 + 6*16
=180 g /mol

!!!!!!!!!!!!!!!!!!!!!.......................................................!!!!!!!!!!!!!!!!!!!!!!!
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Molar mass of glucose (C6H12O6) = 6*C + 12*H + 6*O = 6*12 + 12*1 + 6*16 =180 g /mol
 
Thanks and regards,
Kushagra